OpenStudy (anonymous):

integrate (x+2)/(x+1) dx ????

OpenStudy (anonymous):

I feel like you're supposed to integrate by parts, but I don't know how that would work....?

OpenStudy (anonymous):

\[\frac{x+2}{x+1}=1+\frac{1}{x+1}\] the integrate term by term